3. Use Greens theorem to find the area of an ellipse with semi-axes a and b.
Answer #1

Let us consider the ellipse to be with origin as its center, and its a being along X-axis.

Let us consider the first quadrant part of the ellipse, say E, and let A be the area of this part. Then the area of the complete ellipse is 4A (by symmetry of the ellipse). Note that E = {(a cost, bsint) : + € 0,7/41 . Let B be the region bounded by E and the coordinate axes.

Now, the area inside the ellipse is that of B, and is the integral

\iint_BdS=\iint_Bdx~dy

By Green's theorem, we know

1) (14 – N;) der dy = \ \N do + M dy)

Thus, in order to apply Green's theorem to our problem we need M,N so that

M. - Ny = 1

We try the obvious; say

D +9 hip - +4 = N 19

Then

M = + a Ny= 5+ a - M - N = 1

and hence, by Green, we have

\begin{align*}A&=\iint_B(M_x-N_y)~dx~dy\\ &=\int_E(N~dx+M~dy)\\ &=\int_E{-\frac{ay~dx}{a+b}}+{\frac{bx~dy}{a+b}}\end{align*}

Since E = {(a cost, bsint) : + € 0,7/41 , we get dr = -a sint dt and dy=b\cos t~dt. Therefore,

\begin{align*}A&=\int_E{-\frac{ay~dx}{a+b}}+{\frac{bx~dy}{a+b}}\\ &={\frac 1{a+b}}\int_0^{\pi/2}(a^2b\sin^2t+ab^2\cos^2t)~dt\\ &={\frac 1{2(a+b)}}\int_0^{\pi/2}(a^2b(1-\cos2t)+ab^2(1+\cos2t))~dt\\ &={\frac {(a^2b+ab^2)}{2(a+b)}}\int_0^{\pi/2}dt+{\frac {(-a^2b+ab^2)}{2(a+b)}}\int_0^{\pi/2}\cos2t~dt\\ &={\frac {\pi ab}{4}}+{\frac {(-a^2b+ab^2)}{4(a+b)}}\sin2t\mid_0^{\pi/2}\\ &={\frac {\pi ab}{4}}\end{align*}

Therefore, area of the ellipse is

\begin{align*}4A&={\frac {4\pi ab}{4}}\\ &=\pi ab\end{align*}